OpenStudy (anonymous):

HELP!!! I will give a medal!! Part A: Use the properties of exponents to explain why 81/3 is called the cube root of 8. (5 points) Part B: The length of a rectangle is 3 units and its width is √3 unit. Is the area of the rectangle rational or irrational? Justify your answer. (5 points)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

Anyone?

Nnesha (nnesha):

are u there ???

OpenStudy (whpalmer4):

@katienicole sorry, have been away from OS for a while Use properties of exponents to explain why \(8^{1/3} = \sqrt[3]{8}\) \[x^a * x^b = x^{a+b}\] This means that if \(a=b\) then \[x^a*x^a = x^{a+a} = x^{2a}\] but remember that the square root of a positive number \(x\) is the number \(u\) such that when multiplied by itself gives you the first number: \[u=\sqrt{x}\]\[u*u= u^2 = \sqrt{x}*\sqrt{x} = \sqrt{x*x} =x\] That implies that for square roots, the fractional exponent form would be \(x^{1/2}\) because the sum of the exponents must be \(1\) when we multiply them together, and the only fraction \(a\) which works is \[a+a = 1\]\[2a=1\]\[a=\frac{1}{2}\] So \[\sqrt{x} = x^{1/2}\] because \[x^{1/2}*x^{1/2} = x^{\frac{1}{2} + \frac{1}{2}} = x^1 = x\] You should be able to figure out what this means for cube roots, I hope...

Nnesha (nnesha):

:O :O